rewrite the wquation 4x-10y=11 as a function of x
A) y=-2/5x+11/10
B) y=-2/5x+11
C) y=2/5x-11
D) y=2/5x-11/10

Answer by Nazareth(2)


4x - 10y = 11
Move the -10y to the right of equal sign.
That is add 6y on both sides.
You wil be left with 4x = 11 + 6y
Eliminate the number 11, therefore take away 11 on both side.
You have:
10y = 4x - 11
Divide both side by 10
Answer:
y = 4/10 x - 11/10
Simplify the answer
to have y on one side and x on the other side of =
Therefore y = 2/5 x -11/10

























































































You must tick D for the right answer